### v2

@@ 7.1 Generic vector integrals of motion and their parity properties
Let us note that operators $Q^{(8)},Q^{(7)}$ and $Q^{(6)}$ are equivalent to $Q^{(1)},Q^{(3)}$ and $Q^{(2)}$ respectively up to transformation (24) and so it is sufficient to consider only the potential integrals of motion represented in (139)-(141). The generic vector integral of motion is a linear combination of operators (139)-(143). However, such linear combination can be a priori simplified.
Operators (139)-(143) have different invariance properties with respect to transformation $x_{3}\to-x_{3}$. Half part of them including $Q^{(1)},Q^{(4)},Q^{(5)}$ and $Q^{(8)}$ change their signs together with $x_{3}$ while the remaining operators keep their form.
Let us start with operators $Q^{(1)},Q^{(4)},Q^{(5)},Q^{(8)}$ and consider their linear combination:
$Q=c_{1}Q^{(1)}+c_{4}Q^{(4)}+c_{5}Q^{(5)}+c_{8}Q^{(8)}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}.$ (146)
Let the coefficients $c_{1}$ and $c_{8}$ are nontrivial, then up to the dilatation transformation $c_{1}=\pm c_{8}$. Setting $c_{8}=2$ we come to the following version of operator (146):
$Q=\{K_{3},K_{\alpha}\}\pm 2P_{3}P_{\alpha}+c_{4}\{K_{3},P_{\alpha}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}.$ (147)
If only one of the coefficients $c_{1},c_{8}$ is trivial and $c_{4}$ is nontrivial then up to transformation (24) and dilatation we can reduce (146) to the following form:
$Q=\{(K_{3}\pm P_{3}),P_{\alpha}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},$ (148)
while the remaining inequivalent versions of operator (146) are
$Q=P_{3}P_{\alpha}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}$ (149)
and
$c_{4}\{K_{3},P_{a}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}$ (150)
Analogous speculations with the linear combinations of operators $Q_{2},Q_{3},Q_{6}$ and $Q_{7}$ make it possible to specify the following their inequivalent versions:
$\begin{split}&Q=c_{1}\{D,(K_{\alpha}\pm P_{\alpha})\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},(K_{3}\pm P_{3})\}+Q^{(0)}_{\alpha}+c_{3}\varepsilon_{\alpha\nu}\{L_{\nu},P_{3}\}+x_{\alpha}\tilde{\eta},\\ &Q=c_{1}\{D,P_{\alpha}\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},(K_{3}\pm P_{3})\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=c_{1}\{D,(K_{\alpha}\pm P_{\alpha})\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},P_{3}\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=c_{1}\{D,P_{\alpha}\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},P_{\alpha}\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}.\end{split}$ (151)
Concerning the linear combinations of operators (139) - (143) with different parities with respect to the reflection of $x_{3}$ we can note that such combinations of operator $Q^{(1)}_{\alpha}$ with $Q^{(2)}_{\alpha}$ or $Q^{(3)}_{\alpha}$ are equivalent to $Q^{(2)}_{\alpha}$ or $Q^{(3)}_{\alpha}$ up to the shift transformations. The same is true for the bilinear combinations of $Q^{(2)}_{\alpha}$ with $Q^{(5)}_{\alpha}$ and $Q^{(4)}_{\alpha}$ with $Q^{(7)}_{\alpha}$: they can be reduced to $Q^{(5)}_{\alpha}$ and $Q^{(7)}_{\alpha}$ respectively.
We will not discuss the remaining linear combinations since the computing experiments show that they do not correspond to nontrivial integrals of motion.
Thus to classify the PDM systems with cylindric symmetry which admit vector integrals of motion we are supposed to solve the determining equations (14), (16) which correspond to the symmetries represented in equations (139) - (141) and (147) - (151).
@@ 7.2 Selected calculations
$\begin{split}&\partial_{3}M(\tilde{r},x_{3})=\partial_{1}(x_{1}\tilde{N}(\tilde{r},x_{3})),\\ &\partial_{2}\tilde{N}(\tilde{r},x_{3})=0,\ \partial_{1}M(\tilde{r},x_{3})=\partial_{3}(x_{1}\tilde{N}(\tilde{r},x_{3}))\end{split}$ (152)
@@
In contrast with the case of scalar integrals of motion (refer to Table 1) the obtained solution do not include arbitrary functions but only arbitrary parameters $c_{1},c_{2}$ and $c_{3}$. Moreover, parameter $c_{2}$ can be reduced to zero via the shift of $x_{3}$ provided parameter $c_{1}$ is nontrivial. Thus taking into account that at least one of nonzero parameters in parameter in (153) can be reduced to the unity by simultaneous scaling of all independent variables $x_{1},x_{2}$ and $x_{3}$, without loss of generality we can rewrite the latter equation in the following form:
$M=\cos(\lambda)r^{2}+\sin(\lambda)x_{3}+c_{3},\ \tilde{N}=-\cos(\lambda)x_{3}-\sin(\lambda)$ (154)
where parameter $\lambda$ takes two values: $\lambda=0$ or $\lambda=\frac{\pi}{2}$
@@
The corresponding inverse mass and potential are presented in Item 5 of Table 4 where one more integral of motion is indicated also. Thus the considered system is superintegrable.
The next (and the last) example we consider is the system admitting the symmetry presented in (150), i.e., $c_{1}Q^{(4)}_{\alpha}+c_{2}Q^{(5)}_{\alpha}.$ Setting $\alpha=1$ we come to the following determining equations (14):
@@
Solving step by step the determining equations corresponding to the remaining symmetries (141), (142), (143), (147) - (151) we find all inequivalent PDM systems with cylindric symmetry, which admit vector integrals of motion. The obtained results are presented in Tables 2-5.
@@ 8.1 True tensor integrals of motion
The corresponding PDM system appears to be superintegrable, see Item 2 of Table 4.
@@ 9 Superintegrable systems
The next (and final) step of our analysis is the specifications of just these systems. It can be done in two ways.
The first way presupposes solving of the functional equations. Let you have two systems any of which admit one second order integral of motion, and the related mass functions are $M^{(1)}$ and $M^{(2)}$. Then we have to equate to zero their linear combination:
$k_{1}M^{(1)}+k_{2}M^{(2)}=0.$ (169)
and this condition generates a functional equation for the arbitrary elements present in $M^{(1)}$ and $M^{(2)}$.
The second way is to solve the extended systems of the determining equations (14) corresponding to the inequivalent pairs of integrals of motion. In all cases when such extended systems have nontrivial solutions we can obtain the mass functions of the related superintegrable systems.
The same speculations with the inequivalent triplets of integrals of motion and the related mass functions make it possible to specify the inequivalent maximally superintegrable sytems. We will not present the complete set of the corresponding calculations which do not include new elements in comparison with the given above, but restrict ourselves to considering few examples.
9.1 Superintegrable systems admitting scalar integrals of motion
The systems admitting scalar integrals of motion are defined up to arbitrary functions, and to specify such of them which admit as minimum two such integrals it is preferable to use the second approach, i.e. to search for such cases when arbitrary elements $M$ and $N$ satisfy more than one system of the determining equations specified in (76), (78), (79), (84) and (85)-(108).
Let the PDM system includes arbitrary elements satisfying determining equations (85), see Item 5 of Table 1. The corresponding position dependent mass include two arbitrary functions. Let us search for such the arbitrary functions that the related PDM system admits one more second order integral of motion. It is the case when functions $M$ and $N$ satisfies one more system of the determining equations mentioned in the above.
Solving step by step all pairs of equations including (85) and one of of equations (76), (78), (79), (84) and (88)-(108) we obtain the following results. Equation (85) is incompatible with equations (78), i.e, the corresponding systems of equations have no solutions. The remaining pairs of equations are consistent, but only three of them, namely, (85) and (88), (85) and (91), (85) and (83) have good solutions, presented in Items 1, 3 and 3 of Table 3. Solutions for the other pairs of equations are invariant with respect to two parametric Lie groups and so can be ignored as being classified in our previous papers.
Considering the doublets of equations which include (88) we recognize two more inequivalent systems with good solutions. Those are the doublets including systems (95) or (108). One of the mentioned solutions is represented in Item 4 of Table 3. The other solutions are too cumbersome to be placed in the table, and we fix them in equations (171).
@@ 9.1 Superintegrable systems admitting scalar integrals of motion
We will not represent the following steps and the calculation details since any of the related systems includes four or more equations for two unknowns depending on two variables which make them rather simple solvable. The obtained results are represented in Table 3 and equations (172)-(175). Let us however represent the examples of solutions which are missing in the table and mentioned formulae.
An example of solutions which are not ”good” since the related PDM system is invariant with respect to two parametric group is:
$M=\frac{c_{1}}{\tilde{r}^{2}}+\frac{c_{2}x_{3}}{r\tilde{r}^{2}}.$ (170)
Function (170) solves the system of equations (88), (108). However, the related PDM Scrödinger equation is invariant with respect to two parametric Lie group including rotations around the third coordinate axis and dilatations. Such systems were discussed in [39].
One more nice solution looks as:
$M=\frac{c_{1}}{x_{3}^{2}}+\frac{1}{\tilde{r}^{2}}\left(c_{2}+c_{3}\ln\left(\frac{\tilde{r}}{r^{2}}\right)\right).$
The corresponding integrals of motion are $\hat{Q}_{4}$ (56) and $K_{3}^{2}+P_{a}q(\tilde{r},x_{3})P_{a}+\eta$. We will not represent the explicit form of the related functions $q(\tilde{r},x_{3}),\eta$ and $q_{4}(\tilde{r},x_{3}),\tilde{\eta}_{4}$ in (56) since the latter integral of motion can be reduced to $Q_{1}$ via transformation (22). Notice that this transformation keeps operator $\hat{Q}_{4}$ invariant.
@@
$\begin{split}&f=\frac{(r^{4}\pm 1)^{2}x_{3}^{2}\tilde{r}^{2}}{x_{3}^{2}(c_{1}(r^{4}\mp 1)^{2}\tilde{r}^{2}+c_{3}\tilde{r}^{2}(r^{4}\pm 1)+c_{4}r^{2}\tilde{r}^{2})+c_{2}(r^{4}\pm 1\mp 2r^{2}\tilde{r}^{2})\tilde{r}^{2}},\\ &V=\frac{x_{3}^{2}(c_{5}(r^{4}\mp 1)^{2}\tilde{r}^{2}+c_{7}\tilde{r}^{2}(r^{4}\pm 1)+c_{8}r^{2}\tilde{r}^{2})+c_{6}(r^{4}\pm 1\mp 2r^{2}\tilde{r}^{2})\tilde{r}^{2}}{x_{3}^{2}(c_{1}(r^{4}\mp 1)^{2}\tilde{r}^{2}+c_{3}\tilde{r}^{2}(r^{4}\pm 1)+c_{4}r^{2}\tilde{r}^{2})+c_{2}(r^{4}\pm 1\mp 2r^{2}\tilde{r}^{2})\tilde{r}^{2}},\\ &Q_{1}:=K_{3}^{2}\pm P_{3}^{2}-\hat{\eta}-\frac{(1\pm r^{4})(c_{8}x_{3}^{4}+c_{6}((r^{4}\mp 1)^{2}\pm 4x_{3}^{4})+2c_{7}r^{2}x_{3}^{4})}{x_{3}^{2}(r^{4}\mp 1)^{2}}\\ &+\left(\frac{(1\pm r^{4})(c_{4}x_{3}^{4}+c_{2}((r^{4}\mp 1)^{2}\pm 4x_{3}^{4})+2c_{3}r^{2}x_{3}^{4})}{x_{3}^{2}(r^{4}\mp 1)^{2}}\cdot H\right),\end{split}$ (172)
$\begin{split}&Q_{2}=\{K_{3},P_{3}\}+\frac{c_{6}(\tilde{r}^{2}(r^{4}\mp 1)^{2}\pm 2x_{3}^{4}r^{2})-c_{7}(r^{4}\pm 1)x_{3}^{4}-2c_{8}r^{2}x_{3}^{4}}{2x_{3}^{2}(r^{4}\mp 1)^{2}}\\ &-\left(\frac{c_{2}(\tilde{r}^{2}(r^{4}\mp 1)^{2}-2x_{3}^{4}r^{2})-c_{3}(r^{4}\pm 1)x_{3}^{4}-2c_{4}r^{2}x_{3}^{4}}{2x_{3}^{2}(r^{4}\mp 1)^{2}}\cdot H\right);\end{split}$
@@
9.2 Superintegrable systems with combined symmetries
The next class of superintegrable PDM systems which we consider are those ones which admit both the scalar and vector or tensor integrals of motion. Since the systems admitting the non-scalar integral of motion up to the only exception has a rather simple shape defined up to arbitrary parameters, it is relatively easy to solve the functional equations (169) (where $M^{(1)}$ and and $M^{(2)}$ are position dependent masses of the systems with different types of symmetries) in order to select the cases when there exist additional integrals of motion and so the related systems are superintegrable. To verify the completeness of the obtained in this way list of superintegrable systems it is reasonable to verify the consistence of the systems of the determining equations corresponding to the combinations of symmetries missing in this list.
@@ 9.2 Superintegrable systems with combined symmetries
In this way we find the superintegrable systems presented in Table 4 which collects the systems admitting integrals of motion which transform in different way under rotation transformations. More exactly, there are such systems which admit both scalar and vector or tensor integrals of motion. However, the system represented in Item 2 admits only tensorial integrals of motion.
@@
2 $\frac{1}{c_{1}+2c_{2}x_{3}+c_{3}(\tilde{r}^{2}+4x_{3}^{2})}$ $\frac{c_{4}+2c_{5}x_{3}+c_{6}(\tilde{r}^{2}+4x_{3}^{2})}{c_{1}+2c_{2}x_{3}+c_{3}(\tilde{r}^{2}+4x_{3}^{2})}$ $\begin{array}[]{c}P_{3}^{2}+2c_{5}x_{3}+4c_{6}x_{3}^{2}\\ -\left((2c_{2}x_{3}+4c_{3}x_{3}^{2})\cdot H\right),\\ \{P_{3},D\}-P_{n}x_{3}P_{n}+c_{4}x_{3}+c_{5}(2x_{3}^{2}+\tilde{r}^{2})\\ +c_{6}x_{3}(4x_{3}^{2}+3\tilde{r}^{2})-\left((c_{1}x_{3}+c_{2}(2x_{3}^{2}+\tilde{r}^{2})\right.\\ \left.+c_{3}x_{3}(4x_{3}^{2}+3\tilde{r}^{2}))\cdot H\right),\\ P_{1}L_{1}-P_{2}L_{2}\\ -(x_{1}x_{2}(2c_{3}x_{3}+c_{2})\cdot H)+x_{1}x_{2}(2c_{6}x_{3}+c_{5}),\\ \{P_{1},L_{2}\}+\{P_{2},L_{1}\}\\ -((x_{1}^{2}-x_{2}^{2})(2c_{3}x_{3}+c_{2})\cdot H)\\ +(x_{1}^{2}-x_{2}^{2})(2c_{6}x_{3}+c_{5})\end{array}$

### v3

@@ 7.1 Generic vector integrals of motion and their parity properties
Let us note that operators $Q^{(8)},Q^{(7)}$ and $Q^{(6)}$ are equivalent to $Q^{(1)},Q^{(3)}$ and $Q^{(2)}$ respectively up to transformation (30) and so it is sufficient to consider only the potential integrals of motion represented in (126)-(128). The generic vector integral of motion is a linear combination of operators (126)-(130). However, such linear combination can be a priori simplified using the equivalence transformations. More exactly, we are supposed to consider the following inequivalent versions of vector integrals of motion (see [37])
$\begin{split}&Q=\{K_{3},K_{\alpha}\}\pm 2P_{3}P_{\alpha}+c_{4}\{K_{3},P_{\alpha}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=\{(K_{3}\pm P_{3}),P_{\alpha}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=\{(K_{3}\pm P_{3}),P_{\alpha}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=P_{3}P_{\alpha}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=c_{4}\{K_{3},P_{a}\}+c_{5}\varepsilon_{\alpha\nu}DL_{\nu}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}\end{split}$ (134)
$\begin{split}&Q=c_{1}\{D,(K_{\alpha}\pm P_{\alpha})\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},(K_{3}\pm P_{3})\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=c_{1}\{D,P_{\alpha}\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},(K_{3}\pm P_{3})\}+c_{3}\varepsilon_{\alpha\nu}\{L_{\nu},P_{3}\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=c_{1}\{D,(K_{\alpha}\pm P_{\alpha})\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},P_{\alpha}\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta},\\ &Q=c_{1}\{D,P_{\alpha}\}+c_{2}\varepsilon_{\alpha\nu}\{L_{\nu},P_{\alpha}\}+Q^{(0)}_{\alpha}+x_{\alpha}\tilde{\eta}.\end{split}$ (135)
Thus to classify the PDM systems with cylindric symmetry which admit vector integrals of motion we are supposed to solve the determining equations (20), (22) which correspond to the symmetries represented in equations (126) - (128) and (134), (135).
@@ 7.2 Selected calculations
$\begin{split}&\partial_{3}M(\tilde{r},x_{3})=\partial_{1}(x_{1}\tilde{N}(\tilde{r},x_{3})),\\ &\partial_{2}\tilde{N}(\tilde{r},x_{3})=0,\ \partial_{1}M(\tilde{r},x3)=\partial_{3}(x_{1}\tilde{N}(\tilde{r},x_{3}))\end{split}$ (136)
@@
In contrast with the case of scalar integrals of motion (refer to Table 1) the obtained solution do not include arbitrary functions but only arbitrary parameters $c_{1},c_{2}$ and $c_{3}$.
@@
The corresponding inverse mass and potential are presented in Item 6 of Table 4 where one more integral of motion is indicated also. Thus the considered system is superintegrable.
The next (and the last) example we consider is the system admitting the last symmetry presented in (134), i.e., $c_{1}Q^{(4)}_{\alpha}+c_{2}Q^{(5)}_{\alpha}.$ Setting $\alpha=1$ we come to the following determining equations (20):
@@
Solving step by step the determining equations corresponding to the remaining symmetries (128), (129), (130), (134), (135) we find all inequivalent PDM systems with cylindric symmetry, which admit vector integrals of motion. The obtained results are presented in Tables 2-5.
No $f$ $V$ Integrals of motion
@@ 8.1 True tensor integrals of motion
The corresponding PDM system appears to be superintegrable, see Item 3 of Table 4.
@@ 9 Superintegrable systems
The next (and final) step of our analysis is the specifications of just these systems. To do it wee supposed to solve the extended systems of the determining equations (20) corresponding to the inequivalent pairs of integrals of motion.
The same speculations with the inequivalent triplets of integrals of motion and the related mass functions make it possible to specify the inequivalent maximally superintegrable systems.
To specify the PDM systems as minimum two second order scalar integrals of motion we have to search for such cases when arbitrary elements $M$ and $N$ satisfy more than one system of the determining equations specified in (71), (72), (73), (79) and (80)-(103).
Solving step by step all pairs of the mentioned equations we find the superintegrable systems admitting scalar integrals of motion. The obtained results are presented in Table 3 and the following equations (152)-(153).
The next class of superintegrable PDM systems which we consider are those ones which admit both the scalar and vector or tensor integrals of motion. To specify such systems we have to verify the consistence of the systems of the determining equations corresponding to the combinations of symmetries. In this way we obtain the results presented in Table 4.
@@
$\begin{split}&f=\frac{(r^{4}\pm 1)^{2}x_{3}^{2}\tilde{r}^{2}}{x_{3}^{2}(c_{1}(r^{4}\mp 1)^{2}\tilde{r}^{2}+c_{3}\tilde{r}^{2}(r^{4}\pm 1)+c_{4}r^{2}\tilde{r}^{2})+c_{2}(r^{4}\pm 1\mp 2r^{2}\tilde{r}^{2})\tilde{r}^{2}},\\ &V=\frac{x_{3}^{2}(c_{5}(r^{4}\mp 1)^{2}\tilde{r}^{2}+c_{7}\tilde{r}^{2}(r^{4}\pm 1)+c_{8}r^{2}\tilde{r}^{2})+c_{6}(r^{4}\pm 1\mp 2r^{2}\tilde{r}^{2})\tilde{r}^{2}}{x_{3}^{2}(c_{1}(r^{4}\mp 1)^{2}\tilde{r}^{2}+c_{3}\tilde{r}^{2}(r^{4}\pm 1)+c_{4}r^{2}\tilde{r}^{2})+c_{2}(r^{4}\pm 1\mp 2r^{2}\tilde{r}^{2})\tilde{r}^{2}},\\ &Q_{1}:=K_{3}^{2}\pm P_{3}^{2}-\hat{\eta}-\frac{(1\pm r^{4})(c_{8}x_{3}^{4}+c_{6}((r^{4}\mp 1)^{2}\pm 4x_{3}^{4})+2c_{7}r^{2}x_{3}^{4})}{x_{3}^{2}(r^{4}\mp 1)^{2}}\\ &+\left(\frac{(1\pm r^{4})(c_{4}x_{3}^{4}+c_{2}((r^{4}\mp 1)^{2}\pm 4x_{3}^{4})+2c_{3}r^{2}x_{3}^{4})}{x_{3}^{2}(r^{4}\mp 1)^{2}}\cdot H\right),\\ &Q_{2}=\{K_{3},P_{3}\}+\frac{c_{6}(\tilde{r}^{2}(r^{4}\mp 1)^{2}\pm 2x_{3}^{4}r^{2})-c_{7}(r^{4}\pm 1)x_{3}^{4}-2c_{8}r^{2}x_{3}^{4}}{2x_{3}^{2}(r^{4}\mp 1)^{2}}\\ &-\left(\frac{c_{2}(\tilde{r}^{2}(r^{4}\mp 1)^{2}-2x_{3}^{4}r^{2})-c_{3}(r^{4}\pm 1)x_{3}^{4}-2c_{4}r^{2}x_{3}^{4}}{2x_{3}^{2}(r^{4}\mp 1)^{2}}\cdot H\right);\end{split}$ (154)
@@
10 Discussion
@@
2 $\frac{1}{c_{1}+2c_{2}x_{3}+c_{3}(\tilde{r}^{2}+4x_{3}^{2})}$ $\frac{c_{4}+2c_{5}x_{3}+c_{6}(\tilde{r}^{2}+4x_{3}^{2})}{c_{1}+2c_{2}x_{3}+c_{3}(\tilde{r}^{2}+4x_{3}^{2})}$ $\begin{array}[]{c}P_{3}^{2}+2c_{5}x_{3}+4c_{6}x_{3}^{2}\\ -\left((2c_{2}x_{3}+4c_{3}x_{3}^{2})\cdot H\right),\\ \{P_{3},D\}-P_{n}x_{3}P_{n}+c_{4}x_{3}+\\ c_{5}(2x_{3}^{2}+\tilde{r}^{2})+c_{6}x_{3}(4x_{3}^{2}+3\tilde{r}^{2})\\ -\left((c_{1}x_{3}+c_{2}(2x_{3}^{2}+\tilde{r}^{2})\right.\\ \left.+c_{3}x_{3}(4x_{3}^{2}+3\tilde{r}^{2}))\cdot H\right),\\ P_{1}L_{1}-P_{2}L_{2}\\ -(x_{1}x_{2}(2c_{3}x_{3}+c_{2})\cdot H)\\ +x_{1}x_{2}(2c_{6}x_{3}+c_{5}),\\ \{P_{1},L_{2}\}+\{P_{2},L_{1}\}\\ -((x_{1}^{2}-x_{2}^{2})(2c_{3}x_{3}+c_{2})\cdot H)\\ +(x_{1}^{2}-x_{2}^{2})(2c_{6}x_{3}+c_{5})\end{array}$